OpenStudy (anonymous):

decimal 38 convert to fraction

OpenStudy (anonymous):

.38? .__ tenth place .0__ hundredth place .00__thousandth place .1= 1/10 .01=1/100 .001=1/1000 see the pattern?

OpenStudy (anonymous):

yes thanks